If Z=a+bi/c+di where a, b, c and d are real quantities, show that (a) if z is real then a/b=c/d and (b) if z is entirely imaginary then a/b=-d/c


1
Expert's answer
2021-10-21T14:31:51-0400
"z=\dfrac{a+bi}{c+di}=\dfrac{a+bi}{c+di}\cdot\dfrac{c-di}{c-di}"

"=\dfrac{(ac+bd)+(-ad+bc)i}{c^2+d^2}"

(a) if "z" is real


"-ad+bc=0=>ad=bc=>a\/b=c\/d"

(b) if "z" is entirely imaginary 


"ac+bd=0=>ac=-bd=>a\/b=-d\/c"